One leg of a right triangle is 14 meters longers than the other leg. The hypotenuse is 26 meters long. Find the length of each leg.
Answer by checkley77(12844)

X^2+(X+14)^2=26^2
X^2+X^2+28X+196=676
2X^2+28X+196-676=0
2X^2+28X-480=0
2(X^2+14X-240)=0
2(X+24)(X-10)=0
X-10=0
X=10 M. ANS. FOR THE SHORTER LEG.
10=14=24 M. ANS. FOR THE LONGER LEG.
PROOF:
10^2+24^2=26^2
100+576=676
676=676